(Variable Separable). Determine the solution of the differential equation.

a.) "\alpha" d"\beta" + "\beta"d"\alpha" + "\alpha""\beta"(3d"\alpha" + d"\beta") =0


b.) (xy + x)dx = (x2y2 + x2 + y2 +1)dy


c.) dx = t(1 + t2) sec2x dt


1
Expert's answer
2021-10-25T14:47:00-0400

a)


"\alpha(1+\beta)d\beta=-\beta(1+3\alpha)d\alpha"

"\dfrac{1+\beta}{\beta}d\beta=-\dfrac{1+3\alpha}{\alpha}d\alpha"

Integrate


"\int \dfrac{1+\beta}{\beta}d\beta=-\int\dfrac{1+3\alpha}{\alpha}d\alpha"

"\ln(|\beta|)+\beta=-\ln(|\alpha|)-3\alpha+\ln C"

"\beta e^{\beta}=\dfrac{C}{\alpha e^{3\alpha}}"

"\alpha \beta e^{3\alpha+\beta}=C"

b)


"(y^2+1)(x^2+1)dy=x(y+1)dx"

"\dfrac{y^2+1}{y+1}dy=\dfrac{x}{x^2+1}dx"

Integrate


"\int\dfrac{y^2+1}{y+1}dy=\int\dfrac{x}{x^2+1}dx"

"\int\dfrac{y^2+1}{y+1}dy=\int\dfrac{y^2+2y+1-2y-2+2}{y+1}dy"

"=\int(y+1)dy-2\int dy+2\int \dfrac{dy}{y+1}"

"=\dfrac{y^2}{2}+y-2y+2\ln(|y|)+C_1"

"\dfrac{y^2}{2}-y+2\ln(|y|)=\dfrac{1}{2}\ln(x^2+1)+\ln C"

c)


"\cos^2x dx = t(1 + t^2) dt"

Integrate


"\int\cos^2x dx =\int t(1 + t^2) dt"

"\dfrac{1}{2}\int(1+\cos(2x))dx=\int(t+t^3)dt"

"\dfrac{1}{2}x+\dfrac{1}{4}\sin(2x)=\dfrac{t^2}{2}+\dfrac{t^4}{4}+C"